OpenStudy (kcruz99):

Which of the following is equal to the expression below? (5^-6)^2 a-5^4 b-1/5^4 c-1/5^12 d- -5^12

OpenStudy (astrophysics):

\[\huge (x^m)^n \implies x^{mn}\]

OpenStudy (kcruz99):

so it would be d?

OpenStudy (chrisdbest):

yes

OpenStudy (chrisdbest):

D is the answer!

OpenStudy (jdoe0001):

\(\bf (5^{-6})^2 \\ \quad \\ (a^n)^m\implies a^{n\cdot m} \\ \quad \\ a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\)

OpenStudy (astrophysics):

Since \[\huge \frac{ 1 }{ x^n } \implies x^{-n}\]

OpenStudy (astrophysics):

\[\huge (5^{-6})^2 = 5^{-12} = \frac{ 1 }{ 5^{12} }\]